16. What is the average fluid velocity in a 5.00mm diameter vein, if a 0.250T field across it creates a 1.0mV Hall voltage?
Answer #1


B -0.257 hen E = ImV -3 = 103 F = gū =DE v B - V = Ē В 3 l.xio 0.25 HX103 m/s andy